OpenStudy (anonymous):

-12x-5=10-7x the answer the book gives is -3. I am not sure how they are getting it.

OpenStudy (anonymous):

12x -5 = 10 - 7x 12x + 7x = 10 + 5 19x = 15/19 x= 19/15 = 1 full and 4/15

OpenStudy (anonymous):

12x - 5 = 10 - 7x 12x + 7x = 10 + 5 19x = 15 19/15 = x

OpenStudy (anonymous):

-12x-5=10-7x try this

myininaya (myininaya):

hmmm.... How is 15/19 equal 19/15?

myininaya (myininaya):

The thing is is that it isn't. Also you might want to get the asker involved instead of just providing all the works. Thanks.

myininaya (myininaya):

The book is wrong if that is the right problem you have written.

OpenStudy (anonymous):

Myininaya is right. If the book has given you the answer -3, it is wrong. Talk to your teacher about the error.

OpenStudy (anonymous):

ill repost this seeing how it was completely ignored -12x-5=10-7x try this missing a negative sign

OpenStudy (anonymous):

you are correct. the correct equation is -12x-5=10-7x. The book gives and answer of -3, I still do not understand how they got that answer.

OpenStudy (anonymous):

alright then, the objective here is to isolate x so its alone on one side we want to move the terms with x over to one side, in otherwords, get rid of the x term from one side this can be accomplished if we add 12x to both sides of the equation what happens is +12x -12x =0, so the x term cancels out on the left side then you just combine and simplify like terms on the right -12x-5=10-7x -12x+12x-5=10-7x+12x -5=10 (-7x+12x)

OpenStudy (anonymous):

now we want to move all the constant terms over to the other side of the variable terms, @mdgracia any ideas how you would do this?

OpenStudy (anonymous):

subtract the -5 from each side?

OpenStudy (anonymous):

try again if we subtracted 5 from both sides -5=10 -7x+12x -5-5=10-5 -7x+12x we would get -10=5 -7x+12x

OpenStudy (anonymous):

we want to remove the constant from the side with the variable what constant is on the side of the variables?

OpenStudy (anonymous):

so bring the 10 over to the left side.

OpenStudy (anonymous):

yes and how would you do that?

OpenStudy (anonymous):

add 10 to both sides?

OpenStudy (anonymous):

close but no we have a 10 on the right side if we added 10 to both sides, 10+10=20 if you dont add, you ___?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

ok, -5=10 -7x+12x -5=10+5x i want you to subtract 10 from both sides,then write out the result

OpenStudy (anonymous):

then I have -15=5x, divide by the 5x and get -3. right?

OpenStudy (anonymous):

no, you're solving for x so you want isolate x so its alone without any multiples in other words, you want to turn 5x into x what do you have to do? If you divided by 5x on both sides ( you always need to do it to both sides otherwise the equation will not be balanced) -15/5x=5x/5x \[\frac{-3}{x}=1\] does look right does it?

OpenStudy (anonymous):

Yeah, I did it on both sides. Thanks for walking me through it. I hate math and I am going back and trying to fill in the gaps in my knowledge so I can get better at it.

OpenStudy (anonymous):

the last step would just be dividing 5 from both sides of the equation not dividing 5x you will then get x=-3 the equation "x=-3" is your answer, not "-3"

OpenStudy (anonymous):

That is what I wrote. It is like learning a new language for me. Thanks for the help.